The magnetic flux through a coil of wire containing two loops changes at a constant rate from -58 Wb to 38 Wb in 0.34 s. What is the emf induced in the coil?


1
Expert's answer
2022-04-17T17:07:21-0400

The emf induced in the coil is given by the Faraday's law:


"\mathcal{E} = -\dfrac{\Delta\Phi_B}{\Delta t}"

where "\Delta\Phi_B = 38 Wb - (-58 Wb) = 66Wb" is the change in magnetic flux, and "\Delta t = 0.34s" is the time elapsed. Thus, obtain:


"\mathcal{E} = -\dfrac{66Wb}{0.34s} \approx 1.9\times 10^2V"

Answer. "1.9\times 10^2V".